OpenStudy (anonymous):

When 2 less than a number n is increased by 70%, the result is 8.5.

OpenStudy (anonymous):

I know its n-2 __=8.5

OpenStudy (anonymous):

i dont know the rest

OpenStudy (anonymous):

70/100

OpenStudy (anonymous):

(n-2)*1.75=8.5

OpenStudy (anonymous):

isnt it .70

OpenStudy (anonymous):

(n-2)*1.70=8.5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then how do u solve for n

OpenStudy (anonymous):

im sorry i got moved to honors 8th grade math and im new to this

OpenStudy (anonymous):

n=8.5/1.7 + 2

OpenStudy (anonymous):

what?

OpenStudy (ann1987):

1) First write "two less than a number n" That is written as ( n - 2 ) 2) Now write "70% of that amount" That is written as 0.7( n - 2 ) These two amounts add up to 8.5 ( n - 2 ) + 0.7 ( n - 2 ) = 8.5 To solve for n: 1) Get rid of the decimal by multiplying all the terms on both sides by 10 After you have multiplied, you will have this: 10 ( n - 2 ) + 7 ( n - 2 ) = 85 2) Clear the parentheses by distributing the 10 and the 7 After you have distributed, you will get this: 10n - 20 + 7n - 14 = 85 3) Combine like terms After you combine 10n with + 7n, and after you combine - 20 with - 14, you will have: 17n - 34 = 85 4) Add 34 to both sides to isolate the 17n term, like this: 17n = 119 5) Divide both sides by 17 to isolate n n = 7 <-- answer Answer: n = 7

OpenStudy (ann1987):

Check Replace n in the original equation with 7 ( n - 2 ) + 0.7 ( n - 2 ) = 8.5 ( 7 - 2 ) + 0.7 ( 7 - 2 ) should also equal 8.5 5 + 0.7x5 should equal 8.5 5 + 3.5 should equal 8.5 8.5 does equal 8.5 Check